OpenStudy (anonymous):

helloo! this is a geometry question: find the distance between the pair of points whose coordinates are givin. express as a decimal approximation to the nearest hundreth if nessicary. (-8,2) and (-9, 5) PLEASE HELP MEEE! thanks

OpenStudy (anonymous):

Draw a rough picture with those points marked and draw the line between them. Then that is the hypoteneuse of a right triangle whose sides are.....?

OpenStudy (anonymous):

sqrt((5-2)^2+(-9-(-8))^2) = sqrt(9+1) = sqrt(10) = 3.16228

OpenStudy (anonymous):

This comes from pythagorean's theorem : c^2 = a^2 + b^2. c is the distance and a and be are the component vectors of the vector between the points. Or simply: its a triangle.

OpenStudy (anonymous):

there is a formula for finding distance between two points. its \[\sqrt{(x2-x1)^2 + (y2 -y1)}\] the first co ordinates will be the x1 and y1 then the other one is x2 and y2

OpenStudy (anonymous):

the (y2-y1) must be squared also

OpenStudy (anonymous):

Understanding why there is a formula is more important than memorizing the formula.

OpenStudy (anonymous):

right